OpenStudy (anonymous):

find the volume of the solid obtained by revolving the graph of y=1/(xsqrt(x-1)) for 2=

OpenStudy (turingtest):

let's use disk method: here's the pgraph with the radius of the disks labeled: the area of each disk is\[\pi r^2=\pi(\frac1{x\sqrt{x-1}})^2=\frac\pi{x^2(x-1)}\]the integral is from 2 to 3\[V=\pi\int_{2}^{3}\frac1{x^2(x-1)}dx\]integrate with partial fractions I suppose

OpenStudy (turingtest):

here it is with three disks drawn in